nancy invested $6,000 some at 5% and some at 9% annual interest. if her combined annual interest is $400, find the amount in each account.
Answer by flame8855(424)

0.05x + 0.09y = 400
x+y=6000
5x+9y=40000
-5x-5y=-30000
4y=10000
y= 2500 at 5 %
x = 3500 at 9 %